SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (12/18/07)

Attendees: Su Dong, Selina Li, Mark Tibbetts, Rahmat Rahmat, Al Eisner, Haiwen Zhao, Kim Hojeong.

Meeting Information

 Date      :  December 18th, 2007
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • On Saturday, during the colliding beam runs, we had spikes on TSF cell distribution for SL 6 and 7. Pilot did reconfigure but it wasn't cleared. Mike Kelsey confirmed DCH was fine and had no spikes. At this point, beams were gone, so we took two cosmics runs one with COSMICS run type and the other with PHYSICS_LOWLUMI run type. The spikes were gone. Probably whatever caused it was cleared from the second reconfiguration during the run type change. Will keep monitoring. Thanks to all the people helped me over the weekend: Debbie, Selina, Rainer, Su Dong.
  • Today around 2:17pm, pilot paged me via phone, and it didn't go through. The page was about ORC not reconfiguring due to DCZ. Jim powercycled DCZ crates. After operations meeting, I tested the pager, once via phone, once via web. Only the one via web worked. I added warning to the short term instruction.
  • There is a possibility that we will take 3S run earlier than we thought. If it happens earlier than February, we won't have much time to check trigger and filters. I've started to work on it.

EMT

  • All EMT data looks fine.
  • Mark will be away from Sunday til January 11th. While Mark is away if anything happens: Hojeong will do the hot tower masking and Mark will take care of calibration problem from England. Mark's EMT cell phone will be off while he's gone too.

DCT

GLT

Online

L3

DQM

  • All data look good

Trickle injection monitoring

  • Al tagged it (what he presented last week), it's Rainer's choice to use when.

Trigger Configuration

  • Haiwen prepared a summary page on what we know and what we don't know.

AOB

  • We will resume to weekly meeting on January 8th.
  • Su Dong, Selina and Rainer are requested to keep their cell phones close to them during the break, just in case.
  • Happy Holiday everyone!

This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: December 18th, 2007.